Let u satisfy the Laplace equation in a disk Ω = {(x,y) | x2 + y2 < 1} and continuous on ¯Ω. If u(cosθ,sinθ) ≤ sinθ + cos2θ, then show that
u(x,y) ≤ y +x2 −y2, ∀(x,y) ∈ Ω.
1
Expert's answer
2020-09-07T19:00:28-0400

As per the question,

u satisfy the laplace equation

"\Omega" ={(x,y),"x^2+y^2<1"


According to the given condition

u("cos\theta,sin\theta)\le sin\theta+cos2\theta" ...(1)

Let for disk we assume polar coordinates

x"=cos\theta,y=sin\theta"

Since u satisfy the given disk,so u must satisfy the polar coordinates of disk

New equation became

"u(cos\theta,sin\theta)\le sin\theta+cos^2\theta-sin^2\theta"

u(x,y)"\le y+x^2-y^2"

Hence

u(x,y)"\le y+x^2-y^2 \forall(x,y)\isin Q"